Solve for Y(s), the Laplace transform of the solution y(t) to the initial value problem below y* +6y=P - 4. y(0)= 0, y0) =
Answer #1

y + 64 = 1-4 P LÝ ] - 6 LẤY L7 - 4 L1 [s] - 5/0)- Yoj+64f1] -33-५ C+6) {1- 53 4-3 Y(S) Ly= 2 __4 (6) S(S+6) (+6) Y(S) = 4 3